OpenStudy (anonymous):

Find eqn of tangent line to graph of f(x)=4/√x @ (1,4)?? I am stuck on simplifying after plugging in to formula

OpenStudy (amistre64):

when you know a point and slope; how do you form a line?

OpenStudy (amistre64):

(Px,Py), m y-Py = m(x-Px) m = slope = first derivative of f(x)

OpenStudy (anonymous):

so slope is -2x^(-3/2)? i was plugging in to the 'limit as delta x goes to 0' formula

OpenStudy (amistre64):

why would you ever wanna use the limit version? ewww

OpenStudy (amistre64):

f(x)=4x^(-1/2) f'(x) = -2x^(-3/2) is correct; now when x=1 we got our slope

OpenStudy (anonymous):

ok, so f' = -2x/x^(3/2)?, then?

OpenStudy (amistre64):

you got an extra x in that ... prolly a typo

OpenStudy (amistre64):

since we want to know the slope at the point (1,4), plug in for x=1

OpenStudy (anonymous):

i mean -2 for numerator

OpenStudy (amistre64):

1^(...) = 1; so we are left with m=-2

OpenStudy (amistre64):

(Px,Py), m (1 , 4) , -2 y-Py = m(x-Px) y- 4 = -2(x-1) and rearrange to your hearts content

OpenStudy (anonymous):

into y=mx+b ?

OpenStudy (amistre64):

ax+by=c y-Py=m(x-Px) ax/c + by/c = 1 y = mx -mPx + Py y = mx + b any form you want to use ....

OpenStudy (amistre64):

a line is a line is a line regardless of how you want to shape its equation

OpenStudy (anonymous):

i'm sorry the question asked for slope int. form thanks!

OpenStudy (amistre64):

then yes, reshape it into the y=mx+b ;)

OpenStudy (amistre64):

y- 4 = -2(x-1) y- 4 = -2x +2 +4 +4 ------------- y +0 = -2x+6 y = -2x+6